OpenStudy (kise_ryouta):

A sea turtle swims at a speed of 27 km per h. A girl swims 14 dm per s. How much faster does the sea turtle swim than the girl in meters per minute? 1 m = 10 dm 1000 m = 1 km

OpenStudy (kise_ryouta):

do i i put 14/60

OpenStudy (kise_ryouta):

or 1.4/60

OpenStudy (skullpatrol):

First think about what the question is asking for.

OpenStudy (skullpatrol):

How much faster does the sea turtle swim than the girl in meters per minute? What does that^ mean?

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (kise_ryouta):

im thinking

OpenStudy (skullpatrol):

Thinking is good :-)

OpenStudy (kise_ryouta):

hmm compare how fast the turtle is from the girl

OpenStudy (skullpatrol):

Yes, but how are you asked to compare them?

OpenStudy (kise_ryouta):

in meters per minute

OpenStudy (skullpatrol):

Ok, so what does that mean?

OpenStudy (kise_ryouta):

Um to divide

OpenStudy (skullpatrol):

Nope, because they ask: "How much faster does the sea turtle swim than the girl" That^ means they want the difference in speeds, right?

OpenStudy (kise_ryouta):

right

OpenStudy (skullpatrol):

$$\Huge\text{ in meters per minute?}$$

OpenStudy (skullpatrol):

That^ statement tells you to change the units.

OpenStudy (kise_ryouta):

OOOH

OpenStudy (skullpatrol):

They are NOT asking for the ratio of the speeds because a ratio would not have the units of "in meters per minute?"

OpenStudy (kise_ryouta):

oh okay

OpenStudy (instagrammodel):

Good Job, Skull! :-)

OpenStudy (skullpatrol):

So change both given speeds to meters per minute and subtract.

OpenStudy (skullpatrol):

Thanks @InstagramModel

OpenStudy (kise_ryouta):

and how do i change it 0w0

OpenStudy (skullpatrol):

Have you never changed units before?

OpenStudy (kise_ryouta):

you know how you used to know but you just need some memory help yea thats me

OpenStudy (skullpatrol):

$$\Huge 27\dfrac{km}{h}$$ $$\Huge 27\dfrac{km\dfrac{1000m}{1km}}{h}$$

OpenStudy (skullpatrol):

Oops I meant convert the h to minutes.

OpenStudy (kise_ryouta):

i was about to say....anyways its 60 minutes in a hour

OpenStudy (skullpatrol):

$$\Huge 27\dfrac{1000m}{60 min}$$

OpenStudy (kise_ryouta):

27000 meters

OpenStudy (skullpatrol):

What about the minutes?

OpenStudy (kise_ryouta):

its still 60 so 27000m/60min

OpenStudy (skullpatrol):

Good, now simplify.

OpenStudy (kise_ryouta):

2250/5

OpenStudy (phi):

to remember how to convert units I use this idea, which I'll express in an example: say we want to change km to meters then I know I want to "multiply" km by the fraction \( \frac{meters}{km} \) the idea is I want to multiply by a fraction with km in the bottom so that it "divides out" I want meters up top because that is the unit I want to keep thus \[ 27 \ \cancel{km} \cdot \frac{1000 \ meters}{1 \ \cancel{km}} \]

OpenStudy (phi):

if we have to convert (for example) km/hr to m/min then I would do it like this: \[ 27 \frac{\cancel{km}}{hr} \cdot \frac{1000 \ meters}{1 \ \cancel{km}} \] then I want to get rid of the 1/hr so multiply by hr/min \[ 27 \frac{\cancel{km}}{\cancel{hr}} \cdot \frac{1000 \ meters}{1 \ \cancel{km}} \cdot \frac{1 \ \cancel{hr}}{60 \ min}\]

OpenStudy (phi):

to change dm/sec to meters/min \[ \frac{dm}{sec} \cdot \frac{m}{dm} \cdot \frac{sec}{min} \] see how we pick "fractions" that cancel out the dm and cancel out the sec ? of course we need numbers we know 60 sec = 1 min and if we divide both sides by 1 min we get \[ \frac{60 \ sec}{1 \ min} = 1 \] the left side is the fraction we want they tell us 1 m = 10 dm and we want m/dm, so using that equation, divide both sides by 10 dm to get \[ \frac{1 \ m}{10 \ dm} =1 \] and the left side is the conversion fraction we want.

OpenStudy (phi):

putting in numbers: \[ 14 \frac{dm}{sec} \cdot \frac{1 \ m}{10 \ dm} \cdot \frac{60 \ sec}{1 \ min} = 84 \frac{m}{min}\] and similarly \[ 27 \frac{km}{hr} \cdot \frac{1000 \ meters}{1 \ km} \cdot \frac{1 \ hr}{60 \ min}= \frac{27\cdot 1000}{60}\frac{m}{min}= 450 \frac{m}{min} \] now find the difference in the two speeds.

OpenStudy (kise_ryouta):

im lost now

OpenStudy (davejavous):

Following on from what skull patrol was discussing about converting the units: a) What is the speed of the turtle in metres/min? b) What is the speed of the girl in metres/min?